# On the Theory of Bulk Viscosity of Cold Plasmas

### v1

@@ II.2 The Saha equation
$I_{i,a}$ is the $i$-th ionization potential of atom $a$, and $g_{i,a}$ are the statistical weights of the corresponding ions. Here we suppose that the plasma temperature is much higher than the fine energy splitting. For hydrogen, for example, $I_{1,\mathrm{H}}=13.6\,\mathrm{eV}$, $g_{1,\mathrm{H}}=2$ and for the electrons we also have two spin states and $g_{e}=2$.
@@ II.3 Thermodynamic variables
$\gamma_{0}$ $\equiv\gamma(T,\rho)\equiv\dfrac{\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!s}}{\left(\dfrac{\partial p}{\partial\rho}\right)_{\!T}}=\frac{c_{0}^{2}}{c_{\mathrm{N}}^{2}}=\frac{\rho}{p}\left(\frac{\partial p}{\partial\rho}\right)_{\!\!s}=\frac{\rho}{p}\cdot\dfrac{\mathcal{J}}{\mathcal{C}_{v}}$
@@
In the formulae above, the overline denoting thermal equilibrium values have been omitted again. The velocity $c_{0}$ is the sound velocity at low frequency for which during the adiabatic compression of the plasma, the ionization degrees can follow the equilibrium Saha equation. The general formula above has been solved for pure hydrogen plasma [14, Eq. 96]
@@
where $\iota\equiv I/T$, $I=13.6$ eV is the hydrogen ionization potential and $c_{p}\equiv 5/2$ and $c_{v}\equiv 3/2$ are the mono-atomic heat capacities at constant pressure and volume, and $\alpha\equiv\overline{\alpha}$ is the equilibrium ionization degree
@@
where $n_{e}$, $n_{p}$ and $n_{0}$ are the concentrations of electrons, protons and neutral atoms respectively, and we denote the sum of the concentrations of the protons and the neutrals by $n_{\rho}=n_{0}+n_{p}$.
This formula can be generalized for arbitrary low temperature cocktail of alkaline metal and noble gas, for Na-Ne plasma, for example
$\gamma_{0}=\frac{c_{p}[2+\overline{a}_{\mathrm{noble}}(2-\alpha)]+(1-\alpha)\alpha(c_{p}+\iota)^{2}}{c_{v}[2+\overline{a}_{\mathrm{noble}}(2-\alpha)]+(1-\alpha)\alpha\left[(c_{v}+\iota)^{2}+c_{v}\right]},$ (28)
@@
$\overline{a}_{\mathrm{noble}}=\frac{n_{\mathrm{alkaline}}}{n_{\mathrm{noble}}},\quad\overline{\alpha}=\overline{r}_{1,\,\mathrm{alkaline}},\qquad I=I_{1,\,\mathrm{alkaline}}.$ (29)
@@
The central purpose of the present study is to calculate the generalized polytropic index $\hat{\gamma}(\omega)$ describing the entropy production by the ionization-recombination processes parameterized by the complex phase velocity [15, Eq. (81.8)]
@@ III Wannier ridge resonances and Wannier near threshold ionization cross-section
But how does an electron with small energy slightly above the threshold $I$ create the ionization reaction Eq. (35)? The two electrons escape quasi-classically with small energies and again the mirror images are important. This important problem was solved by Wannier in 1953 [16], see also [18, Sec. 147 Near-threshold cross section behavior], who obtained for the cross-section for the reaction Eq. (35)
@@
The numerical values we use $\mathrm{w}\approx 1.18\sim 1$ and $C_{\mathrm{W}}\approx 2.7$ are obtained from the experimental study [19, Fig. 6].
@@ IV.2 Linearization for small density oscillations
$\mathcal{N}_{e}=\frac{n_{e}}{n_{\rho}}=\sum_{a}\overline{a}\sum_{i=1}^{Z_{a}}i\overline{r}_{i,a}(1+\epsilon_{i,a})$
$\dot{\mathcal{N}}_{e}=\frac{n_{e}}{n_{\rho}}=\sum_{a}\overline{a}\sum_{i=1}^{Z_{a}}i\overline{r}_{i,a}\dot{\epsilon}_{i,a}$
@@ IV.3 Numerical method for singly ionized cold plasmas
$\hat{\epsilon}_{0,a}=-\frac{\overline{\alpha}_{a}}{1-\overline{\alpha}_{a}}\hat{\epsilon}_{a},\quad\hat{\epsilon}_{a}\equiv\hat{\epsilon}_{1,a},$ (75)
$\hat{\epsilon}_{a,2}=\hat{\epsilon}_{a,3}=\dots=\hat{\epsilon}_{a,i\,>\,1}=0.$ (76)
@@
$\hat{\epsilon}_{a}=\dfrac{1-\overline{\alpha}_{a}}{1+\overline{\alpha}_{a}\hat{\tilde{\omega}}_{a}}\left\{\dfrac{\iota_{a}}{c_{v}}\hat{\epsilon}_{\rho}-\dfrac{1}{\overline{\mathcal{N}}_{\mathrm{tot}}}\left[\left(1+\dfrac{\iota_{a}}{c_{v}}\right)\sum_{b=1}^{a_{\mathrm{max}}}\iota_{b}\overline{a}_{b}\overline{\alpha}_{b}\hat{\epsilon}_{b}+\left(\iota_{a}+c_{v}+\dfrac{1}{\overline{c}_{e}}\right)\sum_{b=1}^{a_{\mathrm{max}}}\overline{a}_{b}\overline{\alpha}_{b}\hat{\epsilon}_{b}\right]\right\},$ (81)
where
$\hat{\tilde{\omega}}_{a}\equiv\mathsf{D}/f_{a},\qquad f_{a}\equiv f_{1,a}.$ (82)
@@
$\hat{\Gamma}_{a}=\dfrac{1-\overline{\alpha}_{a}}{1+\overline{\alpha}_{a}\hat{\tilde{\omega}}_{a}},\qquad a=1,\,\ 2,\,\dots a_{\mathrm{max}},$ (84)
@@ IV.4 Generalized polytropic index for cold plasmas
Having calculated all necessary deviations from equilibrium, we can further proceed our analysis with the sound waves absorption. For that purpose following [15, Sec. 81, Second viscosity] at real frequency $\omega$ we can introduce complex wave vector $\hat{k}=k^{\prime}+\mathrm{i}k^{\prime\prime}$ and complex phase velocity
@@
$\hat{\gamma}(\omega)$ $\equiv\frac{\hat{\epsilon}_{p}}{\hat{\epsilon}_{\rho}}=1+\frac{\hat{\epsilon}_{{}_{T}}+\overline{c}_{e}\hat{\epsilon}_{e}}{\hat{\epsilon}_{\rho}}$
$=\gamma_{a}-\frac{1}{c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}\hat{\epsilon}_{\rho}}\sum_{a=1}^{a_{\mathrm{max}}}\overline{a}_{a}\overline{\alpha}_{a}\iota_{a}\hat{\epsilon}_{a}$
@@
$\hat{\gamma}(\omega\rightarrow\infty)$ $=\gamma_{a}\equiv 5/3,\quad\gamma^{\prime\prime}(\infty)\rightarrow 0,\quad\gamma^{\prime}(\infty)\rightarrow\gamma_{a},$
@@
V Analysis of the results derived by of the kinetic equation
@@ V.1 Mandelstam-Leontovich and Cole-Cole approximation
The calculated frequency dependence $\hat{\gamma}(\omega)=\gamma^{\prime}+\mathrm{i}\gamma^{\prime\prime}$ deserves to be compared with the Mandelstam-Leontovich (ML) single time-constant approximation cf. [15, Eq. 81.8]
@@
This factor is always high $Q_{a}=k^{\prime}/2k^{\prime\prime}\gg 1$. One can define the real sound velocity [15, Sec. 81, Second viscosity]
@@
The general formula for the wave-vector frequency dependence of the complex wave-vector, cf. Ref. [15, Eq. (81.9-10)]
@@
gives the low frequency limit [15, Eq. (81.9)] for the ML approximation Eq. (107)
@@
and correspondingly high-frequency approximation [15, Eq. (81.10)] ML approximation Eq. (107)
@@
Figure 2: Reciprocal Q-factor of the ML approximation as a function of $x\equiv\omega\tau$ according to Eq. (118) and Eq. (107) $1/Q_{a,\mathrm{max}}\equiv Q_{a,\mathrm{min}}=0.21$ at $x_{\mathrm{max}}=0.8235$. Independently using Eq. (109), we obtain $\omega_{\mathrm{max}}=0.9127$ µrad/s and the corresponding time-constant is determined as $\tau_{\mathrm{max}}=x_{\mathrm{max}}/\omega_{\mathrm{max}}\approx 10.5$ days. In the general case $1/\mathcal{Q}^{\mathrm{(ML)}}$ has a maximum at $x_{\mathrm{max}}=\omega_{\mathrm{max}}\tau_{\mathrm{max}}=\sqrt{\Gamma_{0,\infty}}=c_{0}/c_{\infty}$, cf. [15, Sec. 81]. Here we have plotted only the $1/\mathcal{Q}^{\mathrm{(ML)}}_{a}(\omega)$ from Eq. (118) since the same calculation from Eq. (109) coincides with it as the Cole-Cole plot in Fig. 1
@@
Let us write again the formula for the wave vector with small imaginary part created by bulk viscosity Eq. (114), confer Ref. [15, Eq. (79.6) and Eq. (81.9-10)]
@@
$\zeta_{0}=-\overline{p}\lim_{\omega\rightarrow 0}\frac{\gamma^{\prime\prime}(\omega)}{\omega}.$ (129)
@@
The time-constant $\tau$ from the good working (see Fig. 1) Mandelstam-Leontovich approximation Eq. (107) can be evaluated using [15, Eq. (81.7)]
@@
$\tau_{0}=-\frac{1}{\gamma_{\infty}^{\prime}-\gamma_{0}^{\prime}}\lim_{\omega\rightarrow 0}\frac{\gamma^{\prime\prime}(\omega)}{\omega},$ (134)
@@
Using the so determined $\tau$, one can evaluate [15, Eq. (81.9-10)]
@@
and also Drude frequency dependence of the complex bulk viscosity [15, Eq. (81.6)]
@@
$\tau_{\infty}=\lim_{\omega\rightarrow\infty}\left(\,\frac{1}{\omega}\sqrt{\frac{\zeta_{0}}{\zeta^{\prime}(\omega)}-1}\,\right)\approx\sqrt{\frac{\zeta_{0}}{\omega^{2}\zeta^{\prime}(\omega)}}.$ (140)
@@
VI Qualitatively discussion of the applicability of the results for acoustic heating of plasmas
VI.1 Qualitative considerations

### v2

@@ II.2 The Saha equation
$I_{i,a}$ is the $i$-th ionization potential of atom $a$, and $g_{i,a}$ are the statistical weights of the corresponding ions. Here we suppose that the plasma temperature is much higher than the fine energy splitting. For hydrogen, for example, $I_{1,\mathrm{H}}=13.6\,\mathrm{eV}$, $g_{1,\mathrm{H}}=2$, $g_{0,\mathrm{H}}=1$, and for the electrons we have two spin states and $g_{e}=2$.
@@ II.3 Thermodynamic variables
$\gamma_{0}$ $\equiv\gamma_{0}(T,\rho)\equiv\dfrac{\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!s}}{\left(\dfrac{\partial p}{\partial\rho}\right)_{\!T}}=\frac{c_{0}^{2}}{c_{\mathrm{N}}^{2}}=\frac{\rho}{p}\left(\frac{\partial p}{\partial\rho}\right)_{\!\!s}=\frac{\rho}{p}\cdot\dfrac{\mathcal{J}}{\mathcal{C}_{v}}$
@@
In the formulae above, the overline denoting thermal equilibrium values have been omitted again. The velocity $c_{0}$ is the sound velocity at low frequency for which during the adiabatic compression of the plasma, the ionization degrees can follow the equilibrium Saha equation. The general formula above has been solved for pure hydrogen plasma [14, Eq. 96] and [15, Eq. 17]
@@
where $\iota\equiv I/T$, $I=13.6$ eV is the hydrogen ionization potential and $c_{p}\equiv 5/2$ and $c_{v}\equiv 3/2$ are the mono-atomic heat capacities at constant pressure and volume, and here $\alpha\equiv\overline{\alpha}$ is the equilibrium ionization degree
@@
where $n_{e}$, $n_{p}$ and $n_{0}$ are the concentrations of electrons, protons and neutral atoms respectively, and we denote the sum of the concentrations of the protons and the neutrals by $n_{\rho}=n_{0}+n_{p}$. This formula can be generalized for arbitrary low temperature cocktail of alkaline metal and noble gas [11, Eq. (B1)], (for example Na-Ne plasma)
$\gamma_{0}=\frac{c_{p}[2+\overline{a}_{\mathrm{nbl}}(2-\alpha)]+(1-\alpha)\alpha(c_{p}+\iota)^{2}}{c_{v}[2+\overline{a}_{\mathrm{nbl}}(2\!-\!\alpha)]+(1\!-\!\alpha)\alpha\left[(c_{v}+\iota)^{2}+c_{v}\right]},$ (28)
@@
$\overline{a}_{\mathrm{nbl}}=\frac{n_{\mathrm{nbl}}}{n_{\mathrm{alk}}},\quad\alpha\equiv\overline{\alpha}=\overline{r}_{1,\,\mathrm{alk}},\qquad I=I_{1,\,\mathrm{alk}}.$ (29)
@@
The central purpose of the present study is to calculate the generalized polytropic index $\hat{\gamma}(\omega)$ describing the entropy production by the ionization-recombination processes parameterized by the complex phase velocity [16, Eq. (81.8)]
@@ III Wannier ridge resonances and Wannier near threshold ionization cross-section
But how does an electron with small energy slightly above the threshold $I$ create the ionization reaction Eq. (35)? The two electrons escape quasi-classically with small energies and again the mirror images are important. This important problem was solved by Wannier in 1953 [17], see also [19, Sec. 147 Near-threshold cross section behavior], who obtained for the cross-section for the reaction Eq. (35)
@@
The numerical values we use $\mathrm{w}\approx 1.18\sim 1$ and $C_{\mathrm{W}}\approx 2.7$ are obtained from the experimental study [20, Fig. 6].
@@ IV.2 Linearization for small density oscillations
$\mathcal{N}_{e}$ $=\frac{n_{e}}{n_{\rho}}=\sum_{a}\overline{a}\sum_{i=1}^{Z_{a}}i\overline{r}_{i,a}(1+\epsilon_{i,a}),$
$\dot{\mathcal{N}}_{e}$ $=\frac{n_{e}}{n_{\rho}}=\sum_{a}\overline{a}\sum_{i=1}^{Z_{a}}i\overline{r}_{i,a}\dot{\epsilon}_{i,a},$
@@ IV.3 Numerical method for singly ionized cold plasmas
$\hat{\epsilon}_{a}\equiv\hat{\epsilon}_{1,a},\quad\hat{\epsilon}_{0,a}=-\frac{\overline{\alpha}_{a}}{1-\overline{\alpha}_{a}}\hat{\epsilon}_{a},$ (75)
$\hat{\epsilon}_{a,2}=\hat{\epsilon}_{a,3}=\dots=\hat{\epsilon}_{a,i\,>\,1}=0,$ (76)
$f_{a}\equiv\frac{1}{\tau_{a}}\equiv\beta_{1,a}\overline{n}_{e}.$ (77)
@@
$\hat{\epsilon}_{a}=\dfrac{1-\overline{\alpha}_{a}}{1-\mathrm{i}\omega\tau_{a}\overline{\alpha}_{a}}\left\{\dfrac{\iota_{a}}{c_{v}}\hat{\epsilon}_{\rho}-\dfrac{1}{\overline{\mathcal{N}}_{\mathrm{tot}}}\left[\left(1+\dfrac{\iota_{a}}{c_{v}}\right)\sum_{b=1}^{a_{\mathrm{max}}}\iota_{b}\overline{a}_{b}\overline{\alpha}_{b}\hat{\epsilon}_{b}+\left(\iota_{a}+c_{v}+\dfrac{1}{\overline{c}_{e}}\right)\sum_{b=1}^{a_{\mathrm{max}}}\overline{a}_{b}\overline{\alpha}_{b}\hat{\epsilon}_{b}\right]\right\}.$ (82)
@@
$\hat{\Gamma}_{a}=\dfrac{1-\overline{\alpha}_{a}}{1-\mathrm{i}\overline{\alpha}_{a}\tau_{a}},\qquad a=1,\,\ 2,\,\dots a_{\mathrm{max}},$ (84)
@@ IV.4 Generalized polytropic index for cold plasmas
Having calculated all necessary deviations from equilibrium, we can further proceed our analysis with the sound waves absorption. For that purpose following [16, Sec. 81, Second viscosity] at real frequency $\omega$ we can introduce complex wave vector $\hat{k}=k^{\prime}+\mathrm{i}k^{\prime\prime}$ and complex phase velocity
@@
$\hat{\gamma}(\omega)$ $\equiv\frac{\hat{\epsilon}_{p}}{\hat{\epsilon}_{\rho}}=1+\frac{\hat{\epsilon}_{{}_{T}}+\overline{c}_{e}\hat{\epsilon}_{e}}{\hat{\epsilon}_{\rho}}$ (101)
$=\gamma_{a}-\frac{1}{c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}\hat{\epsilon}_{\rho}}\sum_{a=1}^{a_{\mathrm{max}}}\overline{a}_{a}\overline{\alpha}_{a}\iota_{a}\hat{\epsilon}_{a}$ (102)
@@
$\hat{\gamma}(\omega\rightarrow\infty)=\gamma_{a}\equiv 5/3,\qquad\gamma^{\prime\prime}(\infty)\rightarrow 0,\quad\gamma^{\prime}(\infty)\rightarrow\gamma_{a},$
@@
V Analysis of the results derived by the kinetic equation
@@ V.1 Mandelstam-Leontovich and Cole-Cole approximation
The calculated frequency dependence $\hat{\gamma}(\omega)=\gamma^{\prime}+\mathrm{i}\gamma^{\prime\prime}$ deserves to be compared with the Mandelstam-Leontovich (ML) single time-constant approximation cf. [16, Eq. 81.8]
@@
This factor is always high $Q_{a}=k^{\prime}/2k^{\prime\prime}\gg 1$. One can define the real sound velocity [16, Sec. 81, Second viscosity]
@@
The general formula for the wave-vector frequency dependence of the complex wave-vector, cf. Ref. [16, Eq. (81.9-10)]
@@
gives the low frequency limit [16, Eq. (81.9)] for the ML approximation Eq. (109)
@@
and correspondingly high-frequency approximation [16, Eq. (81.10)] ML approximation Eq. (109)
@@
Figure 2: Reciprocal Q-factor of the ML approximation as a function of $x\equiv\omega\tau$ according to Eq. (120) and Eq. (109) $1/Q_{a,\mathrm{max}}\equiv Q_{a,\mathrm{min}}=0.21$ at $x_{\mathrm{max}}=0.8235$. Independently using Eq. (111), we obtain $\omega_{\mathrm{max}}=0.9127$ µrad/s and the corresponding time-constant is determined as $\tau_{\mathrm{max}}=x_{\mathrm{max}}/\omega_{\mathrm{max}}\approx 10.5$ days. In the general case $1/\mathcal{Q}^{\mathrm{(ML)}}$ has a maximum at $x_{\mathrm{max}}=\omega_{\mathrm{max}}\tau_{\mathrm{max}}=\sqrt{\Gamma_{0,\infty}}=c_{0}/c_{\infty}$, cf. [16, Sec. 81]. Here we have plotted only the $1/\mathcal{Q}^{\mathrm{(ML)}}_{a}(\omega)$ from Eq. (120) since the same calculation from Eq. (111) coincides with it as the Cole-Cole plot in Fig. 1
@@
Let us write again the formula for the wave vector with small imaginary part created by bulk viscosity Eq. (116), confer Ref. [16, Eq. (79.6) and Eq. (81.9-10)]
@@
$\zeta_{0}=-\overline{p}\lim_{\omega\rightarrow 0}\frac{\gamma^{\prime\prime}(\omega)}{\omega}{\color[rgb]{0,0,1}=-\overline{p}\left.\frac{\mathrm{d}\gamma^{\prime\prime}}{\mathrm{d}\omega}\right|_{\omega=0}.}$ (131)
@@
The time-constant $\tau$ from the good working (see Fig. 1) Mandelstam-Leontovich approximation Eq. (109) can be evaluated using [16, Eq. (81.7)]
@@
$\tau_{0}\equiv-\frac{1}{\gamma_{\infty}^{\prime}-\gamma_{0}^{\prime}}\lim_{\omega\rightarrow 0}\frac{\gamma^{\prime\prime}(\omega)}{\omega},$ (136)
@@
Using the so determined $\tau$, one can evaluate [16, Eq. (81.9-10)]
@@
and also Drude frequency dependence of the complex bulk viscosity [16, Eq. (81.6)]
@@
$\tau_{\infty}\equiv\lim_{\omega\rightarrow\infty}\left(\,\frac{1}{\omega}\sqrt{\frac{\zeta_{0}}{\zeta^{\prime}(\omega)}-1}\,\right)\approx\sqrt{\frac{\zeta_{0}}{\omega^{2}\zeta^{\prime}(\omega)}}.$ (142)
@@
VI Hydrogen-helium cocktail analytical solution
In order to illustrate the work of the general scheme, here we analyze the important special case of a cold cocktail of partially ionized hydrogen and neutral helium atoms for the solar chromosphere, i.e.
$\epsilon\equiv\epsilon_{\mathrm{H}}\equiv\epsilon_{1},\qquad\epsilon_{\mathrm{He}}\equiv\epsilon_{2}=0,$ (146)
$\overline{a}_{\mathrm{H}}=1,\qquad\overline{a}_{\mathrm{He}}=0.1\,$, (147)
$1/\tau_{\mathrm{{}_{H}}}=1/\tau_{1}=\beta(T)n_{e}=\beta n_{\rho}\overline{\alpha},$ (148)
$\overline{\mathcal{N}}_{\mathrm{tot}}=1+\overline{\alpha}+\overline{a}_{\mathrm{He}},\qquad\overline{c}_{e}=\frac{\overline{\alpha}}{1+\overline{\alpha}+\overline{a}_{\mathrm{He}}},$ (149)
$\overline{\alpha}(T)\equiv\overline{\alpha}_{\mathrm{H}}=\dfrac{2}{1+\sqrt{4\dfrac{\overline{n}_{\rho}}{n_{q}}+1}},\qquad\overline{\alpha}_{\mathrm{He}}=0.$ (150)
The solution of Saha equation for the hydrogen degree of ionization $\overline{\alpha}(T,\overline{n}_{\rho})$ is given in Ref. [14]. Actually, this is an applicable approximation for the solar chromosphere but the derived formulas are applicable for any alkali-noble cocktail, Na-Ne plasma or arbitrary cold pure element plasma.
In this special case the general Eq. (82) is reduced to a simple linear equation for $\epsilon$ having the solution
$\epsilon=\dfrac{\dfrac{\iota}{c_{v}}\epsilon_{\rho}}{\dfrac{1-\mathrm{i}\omega\tau_{1}\overline{\alpha}}{1-\overline{\alpha}}+\dfrac{\overline{\alpha}}{\overline{\mathcal{N}}_{\mathrm{tot}}}\left[\left(\dfrac{\iota}{c_{v}}+2\right)\!\iota+c_{v}+\dfrac{1}{\overline{c}_{e}}\right]}.$ (151)
The substitution of this solution in the final expression for the generalized polytropic index Eq. (102) gives
$\hat{\gamma}(\omega)$ $=\frac{c_{p}}{c_{v}}-\dfrac{\left(\dfrac{\iota}{c_{v}}\right)^{\!2}(1-\overline{\alpha})\,\overline{\alpha}}{(1-\mathrm{i}\omega\tau_{\mathrm{{}_{H}}}\overline{\alpha})(1+\overline{\alpha}+\overline{a}_{\mathrm{He}})+\left[\left(\dfrac{\iota}{c_{v}}+2\right)\!\iota+c_{v}+\dfrac{1+\overline{\alpha}+\overline{a}_{\mathrm{He}}}{\overline{\alpha}}\right](1-\overline{\alpha})\,\overline{\alpha}}$ (152)
$=\gamma_{a}-\dfrac{(1-\overline{\alpha})\,\overline{\alpha}\,\iota^{2}/c_{v}}{(1-\mathrm{i}\omega\tau_{\mathrm{{}_{H}}}\overline{\alpha})c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}+\left[(c_{v}+\iota)^{2}\overline{\alpha}+c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}\right](1-\overline{\alpha})}$ (153)
$=\frac{c_{p}}{c_{v}}-\dfrac{\left(\dfrac{\iota}{c_{v}}\right)^{\!2}(1-\overline{\alpha})\,\overline{\alpha}}{(1-\mathrm{i}\omega\tau_{\mathrm{{}_{H}}}\overline{\alpha})(1+\overline{\alpha}+\overline{a}_{\mathrm{He}})+\left[\dfrac{(c_{v}+\iota)^{2}}{c_{v}}+\dfrac{1+\overline{\alpha}+\overline{a}_{\mathrm{He}}}{\overline{\alpha}}\right](1-\overline{\alpha})\,\overline{\alpha}}.$ (154)
This solution for the generalized polytropic index for the alkali-noble cocktail Eq. (153) can be rewritten as
$\hat{\gamma}(\omega)=\gamma_{\infty}-\frac{\mathcal{B}}{\mathcal{A}-\mathrm{i}\omega\mathcal{T}},$ (155)
$\mathcal{A}\equiv(2-\overline{\alpha})c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}+(c_{v}+\iota)^{2}(1-\overline{\alpha})\overline{\alpha}$ (156)
$\mathcal{B}=(1-\overline{\alpha})\,\overline{\alpha}\,\iota^{2}/c_{v}$ (157)
$\mathcal{T}\equiv\tau_{\mathrm{{}_{H}}}c_{v}\overline{\alpha}\overline{\mathcal{N}}_{\mathrm{tot}},$ (158)
and in these notations
$\gamma^{\prime\prime}=-\frac{\mathcal{B}\,\omega\mathcal{T}}{\mathcal{A}^{2}+(\omega\mathcal{T})^{2}},\qquad-\left.\frac{\gamma^{\prime\prime}(\omega)}{\omega}\right|_{\omega\rightarrow 0}=\frac{\mathcal{B}\mathcal{T}}{\mathcal{A}^{2}}.$ (159)
It is instructive to note that
$\mathcal{\overline{N}}_{\mathrm{tot}}$ $=1+\overline{a}_{\mathrm{He}}+\overline{\alpha}\equiv\overline{a}_{\mathrm{H}}+\overline{a}_{\mathrm{He}}+\overline{a}_{e}$
$=\sum_{a}^{a_{\mathrm{max}}}\overline{a}_{a}+\overline{a}_{e},\quad\overline{a}_{\mathrm{H}}=1,\quad\overline{a}_{e}=\overline{\alpha},$ (160)
clearly giving the recipe to include more noble gases in the cocktail. Additionally from Eq. (155) we have
$\gamma_{\infty}^{\prime}-\gamma_{0}^{\prime}=\gamma_{a}-\gamma(\omega=0)=\mathcal{B}/\mathcal{A}$ (161)
and the substitution in Eq. (136) gives for the Mandelstam-Leontovich time-constant
$\tau$ $=\frac{\mathcal{T}}{\mathcal{A}}=\frac{\overline{\alpha}c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}}{\mathcal{A}}\tau_{\mathrm{{}_{H}}}$ (162)
$=\frac{c_{v}\overline{\alpha}\overline{\mathcal{N}}_{\mathrm{tot}}}{(2-\overline{\alpha})c_{v}\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)^{2}(1-\overline{\alpha})\overline{\alpha}}\tau_{\mathrm{{}_{H}}}$ (163)
$=\frac{(1\!+\!\overline{a}_{\mathrm{He}}\!+\!\overline{\alpha})\overline{\alpha}c_{v}}{(2\!-\!\overline{\alpha})(1\!+\!\overline{a}_{\mathrm{He}}\!+\!\overline{\alpha})c_{v}+(c_{v}\!+\!\iota)^{2}(1\!-\!\overline{\alpha})\overline{\alpha}}\tau_{\mathrm{{}_{H}}}.$ (164)
This time constant substituted in Eq. (134) gives
$\zeta_{0}$ $=\frac{\overline{p}\mathcal{B}\mathcal{T}}{\mathcal{A}^{2}}=\frac{\overline{p}\mathcal{B}\overline{\alpha}c_{v}\overline{\mathcal{N}}_{\mathrm{tot}}\tau_{\mathrm{{}_{H}}}}{\mathcal{A}^{2}}$ (165)
$=\frac{(1-\overline{\alpha})\,\overline{\alpha}\,\iota^{2}(1+\overline{a}_{\mathrm{He}}+\overline{\alpha})\overline{p}\tau_{\mathrm{{}_{H}}}}{[(2-\overline{\alpha})(1+\overline{a}_{\mathrm{He}}+\overline{\alpha})c_{v}+(c_{v}+\iota)^{2}(1-\overline{\alpha})\overline{\alpha}]^{2}}.$
Finally the important for us high-frequency dispersion-less damping according Eq. (139)
$k^{\prime\prime}(\omega\rightarrow\infty)=\frac{1}{2\tau c_{\infty}}\frac{\gamma_{a}-\gamma_{0}}{\gamma_{a}}=\frac{1}{2\tau c_{\infty}}\frac{\mathcal{B}}{\gamma_{a}\mathcal{A}},$ (166)
$c_{\infty}=\sqrt{\frac{\gamma_{a}\overline{p}}{\overline{\rho}}},\qquad\gamma_{a}=\gamma_{\infty}=\frac{5}{3}.$ (167)
In the general case we have $\hat{k}(\omega)$ expressed by Eq. (107) with $\hat{\gamma}(\omega)$ from Eq. (155). In such a way our general conclusion is that in alkaline-noble gas approximation we have analytical expression for $\zeta_{0}$ and $\tau$ and Mandelstam-Leontovich formula for $\zeta(\omega)$ gives the exact frequency dependence.
VI.1 Special cases analysis
  • For completely ionized hydrogen plasma $\overline{\alpha}=1$ we have $\hat{\gamma}(\omega)=\gamma_{a}$.
  • For atomic hydrogen $\overline{\alpha}=0$ also because for monoatomic gases the bulk viscosity is zero [24].
  • For very high frequencies $\omega\rightarrow\infty$, $\hat{\gamma}(\omega)=\gamma_{a}$ and $\zeta^{\prime}=0$ again because the influence of the ionization-recombination processes in this $\omega\tau\gg 1$ case is negligible.
  • In the opposite case of evanescent frequencies at formally $\omega=0$, we reproduce after some algebra $\hat{\gamma}(\omega=0)=\gamma_{0}$ the thermodynamic result Eq. (28[11, Eq. (B1)], with $\overline{a}_{\mathrm{nbl}}\equiv\overline{a}_{\mathrm{He}}$ and $\alpha\equiv\overline{\alpha}$.
  • For interstellar plasma it is possible $T\ll I$ at significant $(1-\overline{\alpha})\overline{\alpha}$, formally at $(1-\overline{\alpha})\overline{\alpha}\iota^{2}\gg 1$ we obtain $\gamma_{0}\approx\gamma_{a}-1/c_{v}=1$, $c_{0}=c_{\mathrm{N}}$ and $c_{0}^{2}=\overline{p}/\overline{\rho}$. This formula suggested initially by Newton is an important test for the general program for numerical calculation of bulk viscosity of cold plasma.
  • As a function of the frequency $\omega$ Eq. (152) gives a simple algebraic expression linear function divided by another linear function and taking into account the limit cases at low and high frequencies, we conclude that in this singly ionized atomic approximation we arrive at Mandelstam-Leontovich formula Eq. (109) which we have derived in the present study by the microscopic kinetic approach.
  • Finally, knowing the general relation between the bulk viscosity and generalized polytropic index Eq. (129), we can determine the relaxation time $\tau$ of the ML theory. For example, comparing the high frequency behavior Eq. (142) with the low frequency approximation Eq. (136), we again arrive at Eq. (164).
  • Having a reliable approximation for low temperature hydrogen-helium cocktail, we can address some practical for plasma physics problems as the long standing problem of heating of solar chromosphere, for example.
As a concluding remark of this section, it is good to repeat that formulae derived for the hydrogen-helium cocktail of the solar plasma are applicable for laboratory alkaline-noble gas cocktails.
VII Qualitatively discussion of the applicability of the results for acoustic heating of plasmas
VII.1 Qualitative considerations